OpenStudy (anonymous):

I 2x-3 I=5 values of x

OpenStudy (saifoo.khan):

|2x - 3| = 5 Split into two, 2x - 3= 5 and 2x - 3 = -5 Solve both.

OpenStudy (anonymous):

thanks

OpenStudy (saifoo.khan):

Np